OpenStudy (anonymous):

solve for x: 7x^2+4xz-6y-x=8

OpenStudy (amistre64):

solve the quadratic formula after you conform it to ax^2 + bx + c

OpenStudy (amistre64):

7x^2 +4xz -6y -x = 8 (7)x^2 +(4z-1)x + (-6y-8) = 0

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

your welcome; at least thats the simplest way I can see it :)

OpenStudy (anonymous):

O sorry it's 7y^2 instead of 7x^2

OpenStudy (anonymous):

The answer for x is:\[x=\frac{1}{14} \left(1-\left(4 z\pm \sqrt{168 y+16 z^2-8 z+225}\right)\right) \]